A wheel 200mm in diameter coasts to rest from a speed of 800rpm in 600 seconds. Determine the angular acceleration?


1
Expert's answer
2021-03-10T17:19:38-0500
"\alpha=\dfrac{\omega-\omega_0}{t},""\alpha=\dfrac{0-800\ \dfrac{rev}{min}\cdot\dfrac{1\ min}{60\ s}\cdot\dfrac{2\pi\ rad}{1\ rev}}{600\ s}=-0.14\ \dfrac{rad}{s^2}."

The sign minus means that the wheel decelerates.